# Non-local finite-depth circuits for constructing SPT states and quantum cellular automata

### 04 Dec 2023 v3

@@ III.2 Application to compactified higher-dimensional systems
Applying the construction described above then automatically gives a $\mathrm{FDQC}_{k}$ on the supersites realizing $Q$. However, we must be careful to confirm that the depth of the $k$-local circuit and the value of $k$ are both independent of system size in the compactified dimensions. This is clearly true of $V_{R}$, whose definition in Eq. 12 is unaffected by the compactification and hence it is still $k$-local and finite-depth. Note that while the definition of $V_{R}$ depends on which dimensions we choose to compactify, $V_{R}$ is unaffected by whether we actually compactify those dimensions or not. What remains then is to check that the unitary $w$ used to define $W_{1}$ and $W_{2}$ can be realized as an $\mathrm{FDQC}_{k}$. Note that $W_{1}$ and $W_{2}$ each consists of a parallel application of $w$. Then, if $w$ can be realized as an $\mathrm{FDQC}_{k}$, so can $W_{1}$ and $W_{2}$. We show that this is the case by showing that $w$ is in fact a $(d-1)$-dimensional QCA and then using an inductive argument.
We now show that $w$ is locality-preserving as well as transitionally invariant in the compactified dimensions, i.e. it is a $(d-1)$-dimensional QCA. Let $O$ be a local operator. Recall that $w=\bar{v}u$. Let $uOu^{-1}=\sum_{k}A^{k}\otimes B^{k}$ where $A^{k}$ and $B^{k}$ are operators supported only on the $\ell$- and $r$-dimensional Hilbert spaces that come out of $u$, respectively. Using the Schmidt decomposition we can always choose $A^{k}$ ($B^{k}$) to come from a linearly independent set of operators acting on $\mathbb{C}^{\ell}$ ($\mathbb{C}^{r}$). Let $C^{k}=\bar{v}A^{k}\bar{v}^{-1}$ and $D^{k}=\bar{v}B^{k}\bar{v}^{-1}$. Note that since $v$ is a unitary map and since $A^{k}$ and $B^{k}$ were linearly independent, $C^{k}$ and $D^{k}$ are linearly independent as well. Then from linearity, we have,
@@
where the bars appearing on $C^{k}$ and $D^{k}$ again indicate spatial reversal since $A^{k}$ and $B^{k}$ have been conjugated by $v$ rather than $\bar{v}$. As is shown in Appendix C, since $QO_{2i,2i+1}Q^{-1}$ is contained around $O$, linear independence ensures that each $\bar{C}^{k}$ and $\bar{D}^{k}$ should be contained around $O$. This in turn shows that $wOw^{-1}$ in Eq. 20 is also contained around $O$, and hence $w$ is also locality-preserving in the $d-1$ compactified dimensions. A similar argument shows that $w$ is translationally invariant in the compactified directions if $Q$ is (see Appendix C for details). In other words, $w$ is a $(d-1)$-dimensional QCA.
@@
We now discuss our result in light of the conjecture stated in Ref. Shirley et al. (2022) which states that QCA realizing Witt-equivalent topological orders should be related by a finite-depth circuit i.e., according to this conjecture, one would expect the existence of a finite-depth circuit relating $\bar{Q}$ and $Q^{-1}$. A QCA $Q$ is naturally associated with the chiral topological order $\mathcal{M}_{Q}$ appearing on, say, the top boundary of the 3D Walker-Wang model achieved by truncating the model prepared by QCA $Q$ on a torus. The bottom boundary will support a topological order of the opposite chirality such that the slab as a whole gives a Witt-trivial topological order, i.e. one that admits a gapped boundary to vacuum. Then, if we take $\bar{Q}$, which exchanges the top and bottom boundaries relative to $\mathcal{M}_{Q}$, we have $\mathcal{M}_{\bar{Q}}\sim(\mathcal{M}_{Q})^{-1}$ where $\sim$ indicates the equivalence in terms of boundary topological order. Similarly, if we consider $Q^{-1}$, we have $\mathcal{M}_{Q^{-1}}\sim(\mathcal{M}_{Q})^{-1}$, which gives $\mathcal{M}_{\bar{Q}}\sim\mathcal{M}_{Q^{-1}}$. Our $k$-local unitary $w$, that relates $\bar{Q}$ and $Q^{-1}$ is a circuit whose depth is a constant function of the dimension orthogonal to the truncated boundaries of the Walker-Wang model. Since the choice of the truncation direction is arbitrary, this implies we can write down a different choice of $w$ for each direction such that its depth is constant along that direction. This is the only requirement for $\bar{Q}$ and $Q^{-1}$ to lead to Witt-equivalent boundary topological orders at the top and bottom boundaries of the associated slabs respectively. Hence, this leads to a refinement of the above-mentioned conjecture of Ref. Shirley et al. (2022) to possibly include shifts along all but one direction in the relation of $\bar{Q}$ and $Q^{-1}$ such that they result in Witt-equivalent boundary topological order after truncation.
@@ III.3 Symmetric QCA and SPT phases
We now turn to the symmetry properties of the $\mathrm{FDQC}_{k}$’s constructed in the previous section. Suppose $Q$ is a symmetric QCA, meaning that it commutes with a global symmetry $U(g)=u(g)^{\otimes N}$ for $g\in G$ (we discuss higher-form symmetries at the end). Note that, in contrast to a symmetric QC, where each gate in the circuit is individually symmetric, here we only require that $Q$ as a whole is symmetric, $[Q,U(g)]=0$. Let $U_{A}(g)$ ($U_{B}(g)$) denote $U(g)$ acting on the $A$ ($B$) subsystem as defined in the previous section. $U_{A}(g)U_{B}(g)$ clearly commutes with the swap operators $S_{i}$, since it is a tensor product of the same operator $u(g)$ on every site. Given that $Q_{A}$ is assumed to be symmetric, $V_{i}=Q_{A}S_{i}Q_{A}^{-1}$ commutes with $U_{A}(g)U_{B}(g)$ as well, and thus all gates in the circuit $V_{R}$ defined in Eq. 12 commute with $U_{A}(g)U_{B}(g)$ for all $g\in G$. When we view the sites acted on by $V_{R}$ as the ring $\mathcal{R}$, $U_{A}(g)U_{B}(g)$ is just $U(g)$ acting on all sites in the ring, so it is just the global symmetry of the ring which we denote as $U_{\mathcal{R}}(g)$. For $c\leq i\leq d$, $V_{i}$ and $S_{i}$ have trivial support outside $\mathcal{R}$, and hence the fact that they commute with $U_{A}(g)U_{B}(g)$ readily shows that they also commute with $U_{\mathcal{R}}(g)$. Next, it follows from the results of Cirac et al. (2017) that $w$ commutes with $U(g)$ for all $g\in G$, see Appendix C for detail. Therefore, $W_{1}V_{R}W_{2}$ is a symmetric $\mathrm{FDQC}_{k}$ whose gates commute with $U_{\mathcal{R}}(g)$ for all $g\in G$.
@@
This result has implications for SPT phases. Observe that the $\mathrm{FDQC}_{k}$’s used to disentangle the SPT fixed-point states in Sec. II generate the same unitary operators as the non-symmetric FDQCs used to define the states in the first place, which we call SPT entanglers. That is, our constructions didn’t just trivialize the fixed-point states, they achieved the stronger task of trivializing the SPT entanglers themselves. This perspective allows us to apply our results on representing QCA as symmetric $\mathrm{FDQC}_{k}$’s to SPT phases. For example, fixed-point states for a large class of bosonic SPT phases (the “in-cohomology” phases) with global symmetry are given by the cocycle states defined in Ref. Chen et al. (2013). These states are in turn defined by FDQCs which commute with the global symmetry, but they are not symmetric circuits since the individual gates are not symmetric. Now, our construction allows these circuits to be written as $\mathrm{FDQC}_{k}$’s with symmetric gates. This shows that all in-cohomology SPT phases are trivial in the $k$-local scenario. We demonstrate this idea for the example of the 1D cluster state in Appendix B.2.
@@
Our construction can be applied to some beyond-cohomology SPT phases as well, although these are less well-understood at the Hamiltonian level. One example of a beyond-cohomology SPT phase in 4D with $\mathbb{Z}_{2}$ symmetry was given in Ref. Fidkowski et al. (2020), which also constructed a $\mathbb{Z}_{2}$-symmetric SPT entangler. This entangler is a FDQC, so our construction can be applied to get a symmetric $\mathrm{FDQC}_{k}$. An interesting direction for further work requires extending our results to consider anti-unitary symmetries, i.e. time reversal, which would allow us to address the $k$-local triviality of the 3D beyond-cohomology phase of Ref. Burnell et al. (2014).
@@
Finally, we note that our approach immediately fails for higher-form symmetries. This is because the gates in $V_{R}$ commute only with $U_{A}(g)U_{B}(g)$ and not $U_{A}(g)$ or $U_{B}(g)$ individually. That is, the only symmetry operators which commute with $V_{R}$ are those which act in the same way on the front and back half of $\mathcal{R}$. But the full higher-form symmetry group includes operators that act different on the two halves, as discussed in Sec. II.4. So $V_{R}$ does not commute with the full higher-form symmetry group.
@@ Appendix B Examples of $k$-local circuits for QCA
The 1D shift QCA which translates all sites to the right by one can be represented in the standard form with $d=2$, $\ell=1$, and $r=4$. The $u$ and $v$ operators are defined as,
@@ Appendix E More details on the numerical study of SPT order in monitored random circuits
CNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### 11 Jan 2024 v4

@@ III.2 Application to compactified higher-dimensional systems
Applying the construction described above then automatically gives a $\mathrm{FDQC}_{k}$ on the supersites realizing $Q$. However, we must be careful to confirm that the depth of the $k$-local circuit and the value of $k$ are both independent of system size in the compactified dimensions. This is clearly true of $V_{R}$, whose definition in Eq. 12 is unaffected by the compactification, and hence it is still $k$-local and finite-depth. Note that while the definition of $V_{R}$ depends on which dimensions we choose to compactify, $V_{R}$ is unaffected by whether we actually compactify those dimensions or not. What remains then is to check that the unitary $w$ used to define $W_{1}$ and $W_{2}$ can be realized as an $\mathrm{FDQC}_{k}$. Note that $W_{1}$ and $W_{2}$ each consists of a parallel application of $w$. Then, if $w$ can be realized as an $\mathrm{FDQC}_{k}$, so can $W_{1}$ and $W_{2}$. We show that this is the case by showing that $w$ is in fact a $(d-1)$-dimensional QCA and then using an inductive argument.
We now show that $w$ is locality-preserving as well as transitionally invariant in the compactified dimensions, i.e. it is a $(d-1)$-dimensional QCA. Let $O$ be a local operator. Recall that $w=\bar{v}u$. Let $uOu^{-1}=\sum_{k}A^{k}\otimes B^{k}$ where $A^{k}$ and $B^{k}$ are operators supported only on the $\ell$- and $r$-dimensional Hilbert spaces that come out of $u$, respectively. Using the Schmidt Decomposition we can always choose $A^{k}$ ($B^{k}$) to come from a linearly independent set of operators acting on $\mathbb{C}^{\ell}$ ($\mathbb{C}^{r}$). Let $C^{k}=\bar{v}A^{k}\bar{v}^{-1}$ and $D^{k}=\bar{v}B^{k}\bar{v}^{-1}$. Note that since $v$ is a unitary map and since $A^{k}$ and $B^{k}$ are linearly independent, $C^{k}$ and $D^{k}$ are linearly independent as well. Then from linearity, we have,
@@
In the first equation, we used the Margolus representation of $Q$ and conjugated $O$ by the $u$ operators. The bars appearing on $C^{k}$ and $D^{k}$ again indicate spatial reversal since $A^{k}$ and $B^{k}$ have been conjugated by $v$ rather than $\bar{v}$. As is shown in Appendix C, since $QO_{2i,2i+1}Q^{-1}$ is contained around $O$, linear independence ensures that each $\bar{C}^{k}$ and $\bar{D}^{k}$ should be contained around $O$. This in turn shows that $wOw^{-1}$ in Eq. 20 is also contained around $O$, and hence $w$ is also locality-preserving in the $d-1$ compactified dimensions. A similar argument shows that $w$ is translationally invariant in the compactified directions if $Q$ is (see Lemma 3 in Appendix C for details). In other words, $w$ is a $(d-1)$-dimensional QCA.
@@ III.3 Symmetric QCA and SPT phases
We now turn to the symmetry properties of the $\mathrm{FDQC}_{k}$’s constructed in the previous section. Suppose $Q$ is a symmetric QCA, meaning that it commutes with a global symmetry $U(g)=u(g)^{\otimes N}$ for $g\in G$ (we discuss higher-form symmetries at the end). Note that, in contrast to a symmetric QC, where each gate in the circuit is individually symmetric, here we don’t necessarily have a way to break $Q$ into smaller pieces, so we require only that $Q$ as a whole is symmetric, $[Q,U(g)]=0$. Let $U_{A}(g)$ ($U_{B}(g)$) denote $U(g)$ acting on the $A$ ($B$) subsystem as defined in the previous section. $U_{A}(g)U_{B}(g)$ clearly commutes with the swap operators $S_{i}$, since it is a tensor product of the same operator $u(g)$ on every site. Given that $Q_{A}$ is assumed to be symmetric, $V_{i}=Q_{A}S_{i}Q_{A}^{-1}$ commutes with $U_{A}(g)U_{B}(g)$ as well, and thus all gates in the circuit $V_{R}$ defined in Eq. 12 commute with $U_{A}(g)U_{B}(g)$ for all $g\in G$. When we view the sites acted on by $V_{R}$ as the ring $\mathcal{R}$, $U_{A}(g)U_{B}(g)$ is just $U(g)$ acting on all sites in the ring, so it is just the global symmetry of the ring which we denote as $U_{\mathcal{R}}(g)$. For $c\leq i\leq d$, $V_{i}$ and $S_{i}$ have trivial support outside $\mathcal{R}$, and hence the fact that they commute with $U_{A}(g)U_{B}(g)$ readily shows that they also commute with $U_{\mathcal{R}}(g)$. Next, it follows from the results of Cirac et al. (2017) that $w$ commutes with $U(g)$ for all $g\in G$, see Appendix C for detail. Therefore, $W_{1}V_{R}W_{2}$ is a symmetric $\mathrm{FDQC}_{k}$ whose gates commute with $U_{\mathcal{R}}(g)$ for all $g\in G$.
@@
This result has implications for SPT phases. Observe that the $\mathrm{FDQC}_{k}$’s used to disentangle the SPT fixed-point states in Sec. II generate the same unitary operators as the non-symmetric FDQCs used to define the states in the first place, which we call SPT entanglers. That is, our constructions didn’t just trivialize the fixed-point states, they achieved the stronger task of expressing the SPT entanglers themselves as symmetric $\mathrm{FDQC}_{k}$’s. This perspective allows us to apply our results on representing QCA as symmetric $\mathrm{FDQC}_{k}$’s to SPT phases. For example, fixed-point states for a large class of bosonic SPT phases (the “in-cohomology” phases) with global symmetry are given by the cocycle states defined in Ref. Chen et al. (2013). These states are in turn defined by FDQCs which commute with the global symmetry, but they are not symmetric circuits since the individual gates are not symmetric. Now, our construction allows these circuits to be written as $\mathrm{FDQC}_{k}$’s with symmetric gates. This shows that all in-cohomology SPT phases are trivial in the $k$-local scenario. We demonstrate this idea for the example of the 1D cluster state in Appendix B.2.
@@
Our construction can be applied to some beyond-cohomology SPT phases as well, although these are less well-understood at the Hamiltonian level. One example of a beyond-cohomology SPT phase in 4D with $\mathbb{Z}_{2}$ symmetry was given in Ref. Fidkowski et al. (2020), where a $\mathbb{Z}_{2}$-symmetric SPT entangler was also constructed. This entangler is a FDQC, so our construction can be applied to get a symmetric $\mathrm{FDQC}_{k}$. An interesting direction for further work requires extending our results to consider anti-unitary symmetries, i.e. time reversal, which would allow us to address the $k$-local triviality of the 3D beyond-cohomology phase of Ref. Burnell et al. (2014).
@@
Finally, we note that our approach immediately fails for higher-form symmetries. This is because the gates in $V_{R}$ commute only with $U_{A}(g)U_{B}(g)$ and not $U_{A}(g)$ or $U_{B}(g)$ individually. That is, the only symmetry operators that commute with $V_{R}$ are those that act in the same way on the front and back half of $\mathcal{R}$. But the full higher-form symmetry group includes operators that act differently on the two halves, as discussed in Sec. II.4. So $V_{R}$ does not commute with the full higher-form symmetry group.
@@ Appendix B Examples of $k$-local circuits for QCA
The 1D shift QCA which translates all sites to the right by one, can be represented in the standard form with $d=2$, $\ell=1$, and $r=4$. The $u$ and $v$ operators are defined as,